# Late-time Evolution and Instabilities of Tidal Disruption Disks

### v1

@@ 4.1 Dependence on BH Mass
The early evolution also changes for different $M_{\rm BH}$ values. For $M_{\rm BH}=10^{6}\,M_{\odot}$ or $10^{7}\,M_{\odot}$, the disk accretion rate closely follows the fallback rate, while for $M_{\rm BH}=10^{8}\,M_{\odot}$ the accretion rate is well below the fallback rate except during the high states. The reason for this is that the fallback evolution is slow for high mass BHs, so accretion prevents the disk from building at early times. Whether or not this happens for real TDEs likely depends on the details of the circularization process and when a viscous accretion disk is actually established. It may also depend on whether heating from the fallback stream is important (as discussed in Section 2.4). For all these reasons, we are hesitant to too strongly interpret this early evolution without using a starting point that is more closely set by detailed simulations. Nevertheless, these early uncertainties do not impact the flaring activity that is exhibited later on, which are insensitive to the initial conditions.
Figure 8 shows the same models as in Figure 7, but now on a timescale of 20 years to highlight the longer-term evolution. We see that the waiting time between high states increases with time. This is due to the lower fallback accretion rates, which causes the disk to cycle more slowly between low and high states. The fallback rate continues to drop, but is still high enough to power outburst cycles out to a couple of decades.
@@ 4.2 Mass Ejections During the High State
taken over the high state. This shows that in principle the high state flares can occur out to $\gtrsim 20\,{\rm yrs}$ after the TDE. The typical duration is $\sim 1-2\,{\rm days}$, which is fairly insensitive to the BH mass. The mass ejected during this phase $M_{\rm flare}\sim 10^{-3}-10^{-1}\,M_{\odot}$, with larger BHs ejecting more mass per flare.
@@
4.3 Changes with $M_{*}$ and $\alpha$
@@ 5 Comparisons to Observations
As discussed in Section 1, some of the evidence for long-lasting accretion comes in the form on late-time optical/UV emission and delayed radio flares. We thus compare our disk models with these observations here.
@@ 5.1 Late-time UV Emission
$T_{\rm eff}(r)=\left\{\frac{3GM_{\rm BH}\dot{M}}{8\pi\sigma_{\rm SB}}\left[1-\left(\frac{R_{i}}{r}\right)\right]^{1/2}\right\}^{1/4},$ (18)
where $\sigma_{\rm SB}$ is the Stefan-Boltzmann constant and $R_{i}$ is the inner radius of the disk. Since we are considering BH masses up $10^{8}\,M_{\odot}$, the BH spin must be fairly high to allow the TDE to occur. To be consistent, we assume $R_{i}=R_{s}/2$ (a maximally spinning BH) for our spectral models. Thus there could be strong differences between what we predict and reality at the shortest wavelengths depending on the BH spin. Integrating a Planck function over the entire disk gives the observed flux at a given frequency $\nu$ at a distance $D$ (Frank et al., 2002),
@@
where $\theta$ is the inclination ($\cos\theta=1$ for a face-on disk).
The total luminosity from a disk at a given frequency is calculated by integrating the disk emission over a sphere at radius $D$,
@@
In contrast, an observer calculating the luminosity from an observed flux would typically not take into account the emission pattern and instead simply calculate an isotropic equivalent
@@
Figure 12: Characteristic spectral energy distributions for three different BH masses. In each case, the shaded region represents the range of $\dot{M}$ values possible during the low state (see text for the specific values considered). All models use $M_{*}=M_{\odot}$, $\beta=1.85$, $\alpha=0.1$, and $\cos\theta=1$.
In Figure 12, we consider the range of spectral energy distributions expected for three different BH masses. In each case we use the parameters from the calculations in Figures 7 and 8 ($M_{*}=M_{\odot}$, $\beta=1.85$, and $\alpha=0.1$). We assume the viewing angle is face on, so $\cos\theta=1$. The shaded regions represent the range of accretion rates each BH exhibits while in the low state. For $M_{\rm BH}=10^{6}\,M_{\odot}$, this is $2.6\times 10^{-6}-5.1\times 10^{-4}\,M_{\odot}\,{\rm yr}^{-1}$, for $M_{\rm BH}=10^{7}\,M_{\odot}$, this is $9.8\times 10^{-7}-3.7\times 10^{-4}\,M_{\odot}\,{\rm yr}^{-1}$, and for $M_{\rm BH}=10^{8}\,M_{\odot}$, this is $3.6\times 10^{-7}-3.2\times 10^{-4}\,M_{\odot}\,{\rm yr}^{-1}$. The radius varies as well for each accretion rate, but no more than about a factor of $\sim 3$ for a given BH mass. This shows how the overall disk emission can vary dramatically throughout the low state and how a larger BH mass shifts the distribution to longer wavelengths.
@@
Figure 13: Comparison of the isotropic equivalent luminosity at $150\,{\rm nm}$ for 3 different BH masses with the observations summarized in van Velzen et al. (2019). Detections are shown as filled squares while upper limits are open triangles. Note that this is a combination of observations from 12 different events, and should not be viewed as a light curve (this is further discussed in the text). All models use $M_{*}=M_{\odot}$, $\beta=1.85$, $\alpha=0.1$, and $\cos\theta=1$.
This comparison shows that these models can in principle produce sufficient UV luminosity to explain the observed UV emission. The key factor for achieving this was the inclusion of higher opacities as given by OPAL. One can distinctly see two components during the low state evolution, which is due to the increased opacity from the iron bump. The two upper limits are also consistent with this picture that the late-time UV for any given system may be dimmer by about an order of magnitude as they evolve through the low state. In the future it will be critical to follow the evolution of TDEs in the UV at higher cadence to better understand the duty cycle of this emission.
From Figure 12, it is apparent that the amount of variation in luminosity depends strongly on wavelength. To better highlight this, we compare three different wavebands in Figure 14, where $500\,{\rm nm}$ roughly corresponds to $g$-band and $650\,{\rm nm}$ roughly corresponds to $r$-band. This shows that longer wavelengths generally show less variation. Although there exist some late time $r$-band observations of TDEs (e.g., Nicholl et al., 2024), we caution making direct comparisons to our models because such emission will also have contributions from material at radii slightly beyond $R_{d}$, which is not resolved by our one-zone treatment (it is thus not surprising that our models underpredict the $r$-band luminosity in comparison to these observations). Therefore it will be important to extend our models to 1D in future work to allow more detailed comparisons with observations.
Figure 14: Comparison of how a disk light curve may appear in different wavebands using our $M_{\rm BH}=10^{7}\,M_{\odot}$ model (middle panel from Figure 13). The wavelength of emission is noted in each panel.
@@ 5.2 Radio Flares
As described in Section 4, the super-Eddington phases of the evolution typically ejecta $M_{\rm ej}\sim 10^{-3}-10^{-1}\,M_{\odot}$. The material will have a variety of velocities, roughly scaling with the escape velocity from each radius in the disk. In the outer parts of the disk, $v_{\rm ej}\sim 0.03c$, but near the ISCO it can be $\sim 0.3c$ or somewhat higher. The outflow decreases the amount of material flowing in through each radius (e.g., Blandford & Begelman, 1999)
$\dot{M}_{\rm in}(r)=\left(\frac{r}{R_{d}}\right)^{s}\dot{M},$ (23)
where $s$ is a parameter that can vary from 00 to $1$. Thus, only a small fraction of the material will reach the highest velocities at the inner edge of the disk. Therefore the typical kinetic energy of this ejecta is $\sim 10^{50}\,{\rm erg}$.
Assuming a roughly spherical outflow, this material would be optically thick to the emission from the disk until it has expanded beyond a radius of
$r_{\tau=1}\approx\left(\frac{\kappa M_{\rm flare}}{4\pi}\right)^{1/2}.$ (24)
Up until the ejecta has reached this radius of $r_{\tau=1}\sim 10^{15}\,{\rm cm}$, this material could reprocess the disk emission and increase the optical/UV luminosity. Due to the speed of the ejecta though, it would reach this radius in $\sim{\rm few}\times{\rm days}$, so it might be difficult to catch such a phase.
As the ejecta continues to expand, it will interact with surrounding material with density $n$, sweeping up a mass comparable to its own on at a Sedov-Taylor radius of
@@
where $m_{p}$ is the proton mass. A typical radius would be $r_{\rm TS}\sim 10^{17}-10^{18}\,{\rm cm}$, depending on the value of $n$ which may be higher in the cicumnuclear environment. Thus we would expect radio synchrotron emission on a timescale of $\sim{\rm yrs}$. This is analogous to the radio flare model for neutron star mergers by Nakar & Piran (2011), and we would expect similar luminosities of $\sim 10^{38}\,{\rm erg\,s^{-1}}$ at around $\sim{\rm GHz}$ frequencies. Such a model may have difficulty though producing flares that fall off more abruptly. This could instead require collisions between successive ejections, or even within the same ejection, due to the wide range of outflow velocities expected from different parts of the disk.
Roughly speaking though, when Cendes et al. (2024) fit synchrotron models to real events, they find similar parameters to what we find here. This all suggests that the super-Eddington flares our disk models exhibit could naturally explain the observed radio flares, which we will explore in future work. Since the mass in the ejections and the duration between flares depends strongly on $M_{\rm BH}$, modeling the flares could provide a complementary way to constrain the masses of BHs in TDEs.
@@ 6 Discussion and Conclusions
We explored the evolution of TDE accretion disks on long timescales after the initial stellar destruction. Using a semi-analytic one-zone model, we showed that these disks naturally go into a flaring state due to a thermal instability and continued feeding by fallback accretion. The high super-Eddington state can last for a couple of days, ejecting $\sim 10^{-3}-10^{-1}\,M_{\odot}$ of material at speeds of $\sim 0.03-0.3c$, potentially producing radio flares that can occur up to a decade after the main TDE (e.g., Cendes et al., 2024). This is interspersed with low states where the disk luminosity rises for months to years as it grows in mass. This low state may match the late-time UV luminosities seen years after some TDEs (e.g., van Velzen et al., 2019), and we highlight the importance of the iron-opacity bump in generating sufficiently high luminosities. The UV luminosity can vary by over an order of magnitude in the low state, so we may only be observing events at a time when the disks are especially massive and bright.
@@
We thank Xiaoshan Huang and Daichi Tsuna for many helpful discussions and feedback on a previous draft.
@@ Appendix A Solving for the Disk Evolution
To solve for the time evolution of the disk, Equations (8) and (9) are solved forward in time using explicit time-stepping. The accretion rate can vary by orders of magnitude during its evolution, so the time steps must be similarly flexible. This is done by setting the $i$th time step as
@@
The new radius is then set by using $R_{d}=(J_{d}/M_{d})^{2}/(GM_{\rm BH})$. We then go back to the first bullet point above to start working on the next time step.

### v2

@@ 4.1 Dependence on BH Mass
The early evolution also changes for different $M_{\rm BH}$ values. For $M_{\rm BH}=10^{6}\,M_{\odot}$, the disk accretion rate closely follows the fallback rate, while for $M_{\rm BH}=10^{7}\,M_{\odot}$ or $10^{8}\,M_{\odot}$ the accretion rate is well below the fallback rate except during the high states. The reason for this is that the fallback evolution is slow for high mass BHs, so accretion prevents the disk from building at early times. Whether or not this happens for real TDEs likely depends on the details of the circularization process and when a viscous accretion disk is actually established. It may also depend on whether heating from the fallback stream is important (as discussed in Section 2.4). For all these reasons, we are hesitant to too strongly interpret this early evolution without using a starting point that is more closely set by detailed simulations. Nevertheless, these early uncertainties do not impact the flaring activity that is exhibited later on, which are insensitive to the initial conditions.
Figure 8 shows the same models as in Figure 7, but now on a timescale of 20 years to highlight the longer-term evolution. We see that the waiting time between high states increases with time. This is due to the lower fallback accretion rates, which causes the disk to cycle more slowly between low and high states. The fallback rate continues to drop, but is still high enough to power outburst cycles out to almost a decade.
@@ 4.2 Mass Ejections During the High State
using Equation (10). This shows that in principle the high state flares can occur out to $\lesssim 6\,{\rm yrs}$ after the TDE. The typical duration is $\sim 1-2\,{\rm days}$, which is fairly insensitive to the BH mass. The mass ejected during this phase $M_{\rm flare}\sim 10^{-3}-10^{-1}\,M_{\odot}$, with larger BHs ejecting more mass per flare.
@@
4.3 Changes with $M_{*}$, $\alpha$, and $p$
@@
Finally, in Figure 12, we rerun the simulations with $p=0.2$, corresponding to less mass loss in the high state. The main impact is that the disk grows to larger sizes due to the decreased loss of angular momentum. This in turn increases the viscous time so that the disk evolve thought less outburst cycles.
Figure 12: The same as Figure 8, but using $p=0.2$ instead.
@@ 5 Comparisons to Observations
As discussed in Section 1, some of the evidence for long-lasting accretion comes in the form of late-time optical/UV emission and delayed radio flares. We thus compare our disk models with these observations here.
@@ 5.1 Late-time UV Emission
$T_{\rm eff}(r)=\left\{\frac{3GM_{\rm BH}\dot{M}}{8\pi\sigma_{\rm SB}r^{3}}\left[1-\left(\frac{R_{i}}{r}\right)\right]^{1/2}\right\}^{1/4},$ (22)
where $\sigma_{\rm SB}$ is the Stefan-Boltzmann constant. Radiative transfer effects can complicate the relationship between the observed temperature and $T_{\rm eff}$ (e.g. Done et al., 2012), but since we are only doing simply comparisons at this point we save these details for future work. Since we are considering BH masses up $10^{8}\,M_{\odot}$, the BH spin must be fairly high to allow the TDE to occur. To be consistent, we assume $R_{i}=R_{s}/2$ (a maximally spinning BH) for our spectral models. Thus there could be strong differences between what we predict and reality at the shortest wavelengths depending on the BH spin.
Integrating a Planck function over the entire disk gives the observed flux at a given frequency $\nu$ at a distance $D$ (Frank et al., 2002),
@@
where $\theta$ is the inclination ($\cos\theta=1$ for a face-on disk). The total luminosity from a disk at a given frequency is calculated by integrating the disk emission over a sphere at radius $D$,
@@
In contrast, an observer calculating the luminosity from an observed flux would typically not take into account the disk’s emission pattern and instead simply calculate an isotropic equivalent
@@
Figure 13: Characteristic spectral energy distributions for three different BH masses. In each case, the shaded region represents the range of $\dot{M}$ values possible during the low state (see text for the specific values considered). All models use $M_{*}=M_{\odot}$, $\beta=1.85$, $\alpha=0.1$, $p=0.5$, and $\cos\theta=1$.
In Figure 13, we consider the range of spectral energy distributions expected for three different BH masses. In each case we use the parameters from the calculations in Figures 7 and 8 ($M_{*}=M_{\odot}$, $\beta=1.85$, and $\alpha=0.1$). We assume the viewing angle is face on, so $\cos\theta=1$. The shaded regions represent the range of accretion rates each BH exhibits while in the low state. For $M_{\rm BH}=10^{6}\,M_{\odot}$, this is $9.1\times 10^{-6}-1.5\times 10^{-3}\,M_{\odot}\,{\rm yr}^{-1}$, for $M_{\rm BH}=10^{7}\,M_{\odot}$, this is $3.4\times 10^{-6}-1.0\times 10^{-3}\,M_{\odot}\,{\rm yr}^{-1}$, and for $M_{\rm BH}=10^{8}\,M_{\odot}$, this is $8.8\times 10^{-7}-5.8\times 10^{-4}\,M_{\odot}\,{\rm yr}^{-1}$. The radius varies as well for each accretion rate, but no more than about a factor of $\sim 3$ for a given BH mass. This shows how the overall disk emission can vary dramatically throughout the low state and how a larger BH mass shifts the distribution to longer wavelengths.
@@
Figure 14: Comparison of the isotropic equivalent luminosity at $150\,{\rm nm}$ for 3 different BH masses with the observations summarized in van Velzen et al. (2019). Detections are shown as filled squares while upper limits are open triangles. Note that this is a combination of observations from 12 different events, and should not be viewed as a light curve (this is further discussed in the text). All models use $M_{*}=M_{\odot}$, $\beta=1.85$, $\alpha=0.1$, $p=0.5$, and $\cos\theta=1$.
This comparison shows that these models can in principle produce sufficient UV luminosity to explain the observed UV emission. One can distinctly see two components during the low state evolution, which is due to the increased opacity from the iron bump. The two upper limits are also consistent with this picture that the late-time UV for any given system may be dimmer by about an order of magnitude as they evolve through the low state. In the future it will be critical to follow the evolution of TDEs in the UV at higher cadence to better understand the duty cycle of this emission.
From Figure 13, it is apparent that the amount of variation in luminosity depends strongly on wavelength. To better highlight this, we compare three different wavebands in Figure 15, where $500\,{\rm nm}$ roughly corresponds to $g$-band and $650\,{\rm nm}$ roughly corresponds to $r$-band. This shows that longer wavelengths generally show less variation. This is consistent with the ensemble of light curves presented in (Mummery et al., 2024), which show that the higher cadence g- and r-band light curves generally show variations by a factor of $\sim 3$ rather than being simply flat. This roughly matches the variations we see in Figure 13. Nicholl et al. (2024) also present late time r-band observations of AT2019qiz. The variations appear somewhat smaller, maybe a factor of $\sim 2$, although there are large error bars due to subtracting the host galaxy from the TDE light. To facilitate better comparisons at these longer wavelengths, it will be useful to build one-dimensional models rather than the one-zone models we use here to better follow material at the largest radii where most of this light originates from.
Figure 15: Comparison of how a disk light curve may appear in different wavebands using our $M_{\rm BH}=10^{7}\,M_{\odot}$ model (middle panel from Figure 14). The wavelength of emission is noted in each panel. This demonstrates that the flux variations decrease with longer wavelengths.
@@ 5.2 Radio Flares
As described in Section 4, the super-Eddington phases of the evolution typically ejecta $M_{\rm ej}\sim 10^{-3}-10^{-1}\,M_{\odot}$. The material will have a variety of velocities, roughly scaling with the escape velocity from each radius in the disk. In the outer parts of the disk, $v_{\rm min}\sim 0.03c$, but near the ISCO it can be $v_{\rm max}\sim 0.3c$ or somewhat higher. Using the outflow rate given in Equation (10), this results in a density profile,
$\rho_{\rm out}(v,t)=\frac{p\dot{M}\Delta t}{2\pi(vt)^{3}}\left(\frac{v_{\rm min}}{v}\right)^{2p},$ (27)
where $\Delta t$ is the timescale out of the outburst. Assuming a roughly spherical outflow, the optical depth of this material is
$\tau(t)$ $=$ $\int_{v_{\rm min}t}^{v_{\rm max}t}\kappa\rho(v,t)d(vt)$ (28)
$\approx$ $\frac{p}{4\pi(p+1)}\frac{\kappa\dot{M}\Delta t}{(v_{\rm min}t)^{2}},$
where we assume $v_{\rm min}\ll v_{\rm max}$. Setting $\tau\approx 1$, we can solve for the time when the material becomes optically thin, which is only $\sim 0.5\,{\rm days}$. Thus this material could reprocess the disk emission, and contribute to the optical/UV luminosity, but such a phase may be difficult to catch because it is so short.
Another possibility is that this ejected material generates radio emission. As the ejecta continues to expand, it will interact with surrounding material with density $n$, sweeping up a mass comparable to its own on at a Sedov-Taylor radius of
@@
where $m_{p}$ is the proton mass. A typical radius would be $r_{\rm TS}\sim 10^{17}-10^{18}\,{\rm cm}$, depending on the value of $n$ which may be higher in the cicumnuclear environment. Thus we would expect radio synchrotron emission on a timescale of $\sim{\rm yrs}$. This is analogous to the radio flare model for neutron star mergers by Nakar & Piran (2011), and we would expect similar luminosities of $\sim 10^{38}\,{\rm erg\,s^{-1}}$ at around $\sim{\rm GHz}$ frequencies. Such a model may have difficulty though producing flares that fall off more quickly (e.g., see some of the radio light curves in Alexander et al., 2020). This could instead require collisions between successive ejections when the fast outer material of a later shell catches the slower inner material of an earlier shell. We save a detailed calculation of this process for future work.
Roughly speaking though, when Cendes et al. (2024) fit synchrotron models to real events, they find similar parameters to what we find here. Furthermore, tracing the velocities of the radio emitting regions back in time requires material to be ejected after the main TDE on timescales similar to when we see disk instabilties. This all suggests that the super-Eddington flares our disk models exhibit could naturally explain the observed radio flares, which we will explore in future work. Since the amount of mass in the ejected depends on $M_{\rm BH}$, modeling the flares could provide a complementary way to constrain the BHs in TDEs.
@@ 6 Discussion and Conclusions
We explored the evolution of TDE accretion disks on long timescales after the initial stellar destruction. Using a semi-analytic one-zone model, we showed that these disks naturally go into a flaring state due to a thermal instability and continued feeding by fallback accretion. The high super-Eddington state can last for a couple of days, ejecting $\sim 10^{-3}-10^{-1}\,M_{\odot}$ of material at speeds of $\sim 0.03-0.3c$, potentially producing radio flares that can occur up to a decade after the main TDE (e.g., Cendes et al., 2024). This is interspersed with low states where the disk luminosity rises for months to years as it grows in mass. This low state matches the late-time UV luminosities seen years after some TDEs (e.g., van Velzen et al., 2019). The UV luminosity can vary by over an order of magnitude in the low state, so we may only be observing events at times when the disks are especially massive and bright. This is difficult to assess because current UV observations are fairly sparse. We find that the luminosity in the optical bands varies less, which is similar to the factor of $\sim 3$ variations seen in the late time r- and g-band observations (Mummery et al., 2024).
@@
Our $M_{\rm BH}=10^{6}\,M_{\odot}$ disk model transitions from the high state to the low state for the first time at $\sim 100\,{\rm days}$, depending on the details of the parameters chosen (e.g., see Figure 7). In contrast, the TDE AT2018fyk shows a sudden drop in X-ray emission at $\sim 320$ and $\sim 500\,{\rm days}$ (Wevers et al., 2019, 2023) while AT2021ehb shows a similar drop at $\sim 320\,{\rm days}$ (Yao et al., 2022). This could potentially be matched by more massive BHs. Alternatively, Lu (2022) speculates that these longer timescales could be easier to explain if the disk experiences an outside-in transition to a low accretion state. This process is estimated to roughly take $\sim 1\,{\rm yr}$ (set by when the inner disk reaches the Eddington limit), which is in better agreement with AT2018fyk and AT2021ehb. Addressing whether this occurs would require a one-dimensional treatment rather than the one-zone model we employ here. While this is outside of the scope of this work, such a calculation could be fruitful for better understanding these observations. Another complication that we do not include is the impact of magnetic fields, which can help to stabilize the TDE disks (e.g. Alush & Stone, 2025). Higher cadence observations, especially the the UV with missions like ULTRASAT (Shvartzvald et al., 2024), would be helpful for constraining whether such physical effects are needed.
@@
We thank Xiaoshan Huang, Daichi Tsuna, and Wenbin Lu for helpful discussions and feedback. We also thank the referee for useful comments and suggestions.
@@ Appendix A Solving for the Disk Evolution
To follow the time evolution of the disk, Equations (11) and (12) are solved forward in time using explicit time-stepping. The accretion rate can vary by orders of magnitude during its evolution, so the time steps must be similarly flexible. This is done by setting the $i$th time step as
@@
The new radius is then set by using $R_{d}=(J_{d}/M_{d})^{2}/(GM_{\rm BH})$. We then go back to the first bullet point above to start working on the next time step. Implicit in our approach is that the thermal time of the disk is always much shorter than the viscous time so that energy balance is always maintained. This prevents us from being forced to explicitly time evolve the internal energy of the disk. Our comparisons in Section 3, demonstrate that this assumption results in solutions that match previous work.